Need help with this equation:
Solve the equation if possible if not enter none:
|6-x|+4=18
x= smaller value
x= larger value
thanks

Answer by psbhowmick(878)

|6 - x| + 4 = 18
|6 - x| = 18 - 4
|6 - x| = 14

When x > 6,
-(6 - x) = 14 i.e. x - 6 = 14 i.e. x = 20

When x < 6,
(6 - x) = 14 i.e. 6 - x = 14 i.e. x = -8

Thus, there are two solutions of the equation: x = -8 and x = 20.